# EnScale: Temporally-consistent multivariate generative downscaling via proper scoring rules

### v2

@@ A.4 Results on extrapolation test period
NN-det-extra 1914.9 ± 33.34 4888.1 ± 40.52 4883.9 0.0 29.89 ± 0.28 25.76 ± 0.22 3.5 ± 0.03
EasyUQ-extra 3217.8 ± 40.21 3520.5 ± 29.43 6613.2 6188.4 23.60 ± 0.2 46.98 ± 0.32 10.3 ± 0.02
Analogues-extra 2560.2 ± 36.13 4277.1 ± 44.87 7221.1 5901.3 28.19 ± 0.22 21.02 ± 0.18 1.8 ± 0.02
GAN-extra 2906.5 ± 40.6 3566.5 ± 31.14 6197.6 5265.8 23.43 ± 0.19 17.70 ± 0.15 3.7 ± 0.03
CorrDiff-extra 3466.7 ± 54.06 3100.6 ± 28.9 6243.2 6285.3 21.85 ± 0.19 15.58 ± 0.15 1.5 ± 0.01
EnScale-extra 1936.0 ± 30.34 3371.1 ± 31.24 6930.1 7118.1 23.29 ± 0.19 21.45 ± 0.16 2.9 ± 0.02
EnScale-t-extra 1989.1 ± 29.1 3392.3 ± 30.77 6995.6 7211.4 23.41 ± 0.2 21.89 ± 0.16 2.9 ± 0.02
A.5 Further time series evaluation
Table A.2: Lag-1 autocorrelations (ACF) for each variablea
Method tas pr sfcWind rsds
NN-det 0.01 0.17 0.08 0.11
EasyUQ 0.03 0.08 0.15 0.03
Analogues 0.1 0.12 0.28 0.04
GAN 0.06 0.1 0.21 0.02
CorrDiff 0.01 0.23 0.1 0.09
EnScale 0.08 0.09 0.22 0.06
EnScale-t 0.01 0.04 0.05 0.02
aWe calculate ACFs in each location for both the samples and the RCM data. We average the absolute error (magnitude of the difference of sampled minus target, regardless of the sign) across locations.
Figure A.14: Power spectral densities
Table A.3: RALSD-avg
method tas pr sfcWind rsds
NN-det 0.466 8.598 1.392 4.050
EasyUQ 5.958 7.610 8.501 9.966
Analogues 0.150 0.344 0.153 0.170
GAN 1.999 6.890 3.183 4.066
CorrDiff 0.394 0.951 0.432 0.752
EnScale 0.579 1.456 0.707 1.293
EnScale-t 0.589 1.429 0.654 1.274
Appendix B Details on methods
B.1 Conditional independence assumptions for multi-step approach
@@
B.2 Conditional independence assumptions for time series generation
Formally, our approach in Sec. 2.3 relies on suitable conditional independence assumptions, similar to the above discussion about the downscaling via coarse correction (App. B.1 ). For simplicity, we omit the intermediate step via $\tilde{Z}$ in this description, and just consider the two steps $X_{t}\rightarrow Z_{t}$ and $Z_{t}\rightarrow Y_{t}$. Here, the key approximation is that we assume $Y_{t}$ to not depend on $Y_{t-1}$ when given $Z_{t}$. In other words, only temporal correlations on the coarse scale (in $Z_{t}$) are considered and further dependencies on the local scale (in $Y_{t}$) are neglected.
We test this approximation by comparing the losses in different prediction setups. We predict $Y_{t}$ from $Z_{t}$ only and from $Z_{t},Y_{t-1}$, checking if the assumption $Y_{t}\mathrel{\perp\!\!\!\perp}Y_{t-1}|Z_{t}$ is valid. For simplicity, we predict one variable at a time, using a single dense model as the super-resolution model. In this case, we test precipitation and surface wind. For precipitation, we find no change in loss between the two setups. For surface wind, including $Y_{t-1}$ improves the loss by about 2%. Like above, this suggests that the conditional independence might not be satisfied perfectly, but is a valid approximation that simplifies computations.
More formally, we require $Y_{t}$ to not depend on $Z_{t-1},Y_{t-1},X_{t}$ when given $Z_{t}$. Similarly, $Z_{t}$ should not depend on $Y_{t-1}$ when given $X_{t},Z_{t-1}$. Thus, the assumptions are:
  1. 1.
    $Y_{t}\mathrel{\perp\!\!\!\perp}(X_{t},Z_{t-1},Y_{t-1})|Z_{t}$
  2. 2.
    $Z_{t}\mathrel{\perp\!\!\!\perp}Y_{t-1}|X_{t},Z_{t-1}$
With these, a similar proof to the above (App. B.1) shows that $p_{Y_{t}|X_{t},Y_{t-1}}$ is indeed learned by our model. As before, we drop subscripts and use the corresponding lower-case letters as arguments.
$p(y_{t}|x_{t},y_{t-1})$ $=\int p(y_{t}|x_{t},y_{t-1},z_{t-1},z_{t})p(z_{t-1},z_{t}|x_{t},y_{t-1})d(z_{t-1},z_{t})$
$=\int p(y_{t}|x_{t},y_{t-1},z_{t-1},z_{t})p(z_{t}|x_{t},y_{t-1},z_{t-1})p(z_{t-1}|x_{t},y_{t-1})d(z_{t-1},z_{t})$
$\mathrel{{\mathop{=}\limits^{(1)}}}\int p(y_{t}|z_{t})p(z_{t}|x_{t},y_{t-1},z_{t-1})p(z_{t-1}|x_{t},y_{t-1})d(z_{t-1},z_{t})$
$\mathrel{{\mathop{=}\limits^{(2)}}}\int p(y_{t}|z_{t})p(z_{t}|x_{t},z_{t-1})p(z_{t-1}|x_{t},y_{t-1})d(z_{t-1},z_{t})$
$=\int p(y_{t}|z_{t})p(z_{t}|x_{t},z_{t-1}(y_{t-1}))dz_{t}$
where in the last step we used that, given $Y_{t-1}$, $Z_{t-1}$ is known, and $Z_{t-1}(Y_{t-1})$ denotes the average pooled $Y_{t-1}$. This last row is what the model generates, with one slight simplification: We predict $\hat{Z}_{t}$ from $\hat{Z}_{t-1}$, instead of the average pooled $\hat{Y}_{t-1}$, such that the time-series generation can happen independently of the super-resolution model.
B.3 Batch-wise approximation
The theoretical expected values (e.g. Eq. (2)) are impossible to calculate. However, during training, they can be replaced by averages over mini-batches of data, yielding an unbiased estimate of the Let $(x_{i},y_{i})\,i\in\{1,...,n\}$ be a mini-batch of data, and $\hat{y_{i}},\hat{y_{i}}^{\prime}\sim Q_{Y|X}(.|x_{i})$ two samples from the conditional distribution of the RCM data given the GCM input $x_{i}$. Then the finite-sample approximation of the loss is
@@ B.3 Batch-wise approximation
B.4 Model architecture
B.4.1 Coarse model
The first coarse model is a dense multilayer perceptron (MLP) and its architecture is as in [Shen2024Engression:Regression]. To enable the model to capture difference between GCM-RCM pairs, we integrate both the index for the GCM and RCM considered: We one-hot-encode GCM and RCM index separately and append both to the input. We kept the GCM-index because our work does not consider extrapolation to unseen GCMs, but preliminary tests suggest that the model performance is very similar also without the GCM-encoding. In addition, the MLP is modified to include random noise that allows for stochastic output. As explained above, the generative model needs to take random initialization as input. Instead of just adding this noise to the input at the beginning, we add the randomness at each layer. We concatenate the hidden units with standard Gaussian noise, where the noise is scaled by a learnable parameter. This concatenation is passed into the next layer. Adding noise in multiple hidden layers allows to learn more complex representations of the Gaussian noise. For example, we find that this yields spatially correlated variability in the output.
@@ B.4.1 Coarse model
B.4.2 Super-resolution model
@@
As for the coarse model, we enable learning RCM-specific details by integrating the index for the GCM-RCM pair into the sparse local layers. In the first linear upsampling step, the learnable weights are pair-specific, i.e. the model can learn a different set of weights for each pair. The stochastic refinement does not depend on the GCM-RCM pair and, thus, learned stochasticity is pair-agnostic. We also tested pair-specific stochastic modeling, but found no improvement in the loss.
B.5 Training
@@
B.6 Details on benchmarks
B.6.1 Deterministic neural network
We use a similar model architecture to the dense models in our generative approach (see Sec. B.4), using a fully-connected neural network with alternating linear maps and ReLU activation. However, unlike for the generative models, there is no noise input needed for the deterministic benchmark. The layers have sizes 3612, 500, 500, 500, 1024, 4096, 16384.
@@ B.6.1 Deterministic neural network
B.6.2 EasyUQ
EasyUQ’s predictions are optimal under a monotonicity assumption, in the sense that they minimize the CRPS. The monotonicity assumption describes that the targets $Y$ tend to increase as the regression model’s predictions $\hat{\mu}$ also increase. This is a natural requirement for a useful statistical model. However, if the regression model is deficient, this assumption might be violated. EasyUQ will still yield predictive distributions, but optimality is not guaranteed in this case. For more details on EasyUQ and its mathematical guarantees, see [Walz2024EasyOutput] and [Henzi2021IsotonicRegression].
We implement EasyUQ based on the Python package isodisreg [WalzIsodisreg]. We train EasyUQ on the training data that was also used to train the NN-det, separately for each location. After training, we retrieve the predicted distributions on the test data and sample from them, independently in each location, time point and for each variable. For this, we slightly adjust the default prediction method from the Python package in order to be able to take in a different set of probabilities for each point in time. These samples can then be compared to the samples from EnScale.
B.7 Analogues
@@
B.8 Proof of variance decomposition
@@
We explained above (see App. B.1) that our approach to downscaling via coarse correction assumes $Y\mathrel{\perp\!\!\!\perp}X|Z$. Hence, $\operatorname{Var}^{\iota}(Y|Z,X)=\operatorname{Var}^{\iota}(Y|Z)$ and $E^{\iota}[Y|Z,X]=E^{\iota}[Y|Z]$, which concludes the proof.

### v3

@@ A.7 Results on extrapolation test period
NN-det-extra 1917.0 ± 32.37 4886.2 ± 45.21 4889.5 0.0 29.93 ± 0.27 25.74 ± 0.23 3.5 ± 0.03
EasyUQ-extra 3216.3 ± 42.23 3518.3 ± 31.72 6613.3 6186.4 23.59 ± 0.21 46.95 ± 0.29 10.3 ± 0.03
Analogues-extra 2557.9 ± 38.29 4275.5 ± 47.61 7228.4 5889.6 28.22 ± 0.25 21.04 ± 0.2 1.8 ± 0.02
GAN-extra 2903.0 ± 43.52 3563.3 ± 31.19 6196.1 5271.6 23.46 ± 0.19 17.73 ± 0.14 3.7 ± 0.03
CorrDiff-extra 3457.1 ± 52.99 3100.1 ± 26.18 6240.6 6292.5 21.85 ± 0.21 15.58 ± 0.14 1.5 ± 0.01
CorrDiff-s-extra 3143.7 ± 52.35 3207.3 ± 31.0 6299.2 6166.7 20.95 ± 0.2 15.00 ± 0.14 1.6 ± 0.01
EnScale-extra 1932.3 ± 29.54 3363.9 ± 29.12 6844.5 6945.6 23.24 ± 0.19 20.98 ± 0.14 2.9 ± 0.02
EnScale-t-extra 1947.6 ± 31.29 3373.0 ± 32.02 6885.7 7028.8 23.01 ± 0.2 21.40 ± 0.16 2.8 ± 0.02
Appendix B Extrapolation to unseen GCMs
Table B.1: Scores per GCM-RCM pair. GCM-RCM combinations that were not included in the train data are marked in bold.
tas
RCM GCM MSE ES CRPS
ALADIN63 CNRM-CM5 5.46 184.86 1.29
ALADIN63 HadGEM2-ES 7.31 214.38 1.53
ALADIN63 MPI-ESM-LR 4.89 175.68 1.22
CCLM4-8-17 CNRM-CM5 3.68 154.43 1.09
CCLM4-8-17 CanESM2 6.70 216.06 1.52
CCLM4-8-17 MIROC5 4.09 162.19 1.16
CCLM4-8-17 MPI-ESM-LR 2.53 130.81 0.93
REMO2015 MIROC5 6.25 195.85 1.39
RegCM4-6 CNRM-CM5 5.46 186.48 1.31
RegCM4-6 MPI-ESM-LR 4.82 173.66 1.24
pr
RCM GCM MSE ES CRPS
ALADIN63 CNRM-CM5 40.04 502.29 2.33
ALADIN63 HadGEM2-ES 40.74 517.32 2.26
ALADIN63 MPI-ESM-LR 29.03 413.21 1.86
CCLM4-8-17 CNRM-CM5 31.01 432.09 1.85
CCLM4-8-17 CanESM2 20.21 326.74 1.37
CCLM4-8-17 MIROC5 32.92 441.67 1.73
CCLM4-8-17 MPI-ESM-LR 26.21 394.10 1.69
REMO2015 MIROC5 52.25 557.67 2.20
RegCM4-6 CNRM-CM5 21.93 366.25 1.79
RegCM4-6 MPI-ESM-LR 21.49 356.76 1.73
sfcWind
RCM GCM MSE ES CRPS
ALADIN63 CNRM-CM5 2.82 140.44 0.87
ALADIN63 HadGEM2-ES 3.55 159.83 1.00
ALADIN63 MPI-ESM-LR 2.25 124.98 0.78
CCLM4-8-17 CNRM-CM5 2.55 133.71 0.84
CCLM4-8-17 CanESM2 3.06 147.64 0.94
CCLM4-8-17 MIROC5 2.57 133.34 0.84
CCLM4-8-17 MPI-ESM-LR 1.99 116.87 0.75
REMO2015 MIROC5 3.24 149.61 0.97
RegCM4-6 CNRM-CM5 3.59 158.37 1.04
RegCM4-6 MPI-ESM-LR 3.39 152.52 1.01
rsds
RCM GCM MSE ES CRPS
ALADIN63 CNRM-CM5 2141.77 3659.38 25.04
ALADIN63 HadGEM2-ES 1990.96 3567.66 24.57
ALADIN63 MPI-ESM-LR 1860.73 3395.63 23.05
CCLM4-8-17 CNRM-CM5 2121.95 3478.42 24.30
CCLM4-8-17 CanESM2 2136.50 3537.87 24.29
CCLM4-8-17 MIROC5 2135.70 3496.39 24.28
CCLM4-8-17 MPI-ESM-LR 1895.41 3327.91 22.88
REMO2015 MIROC5 3177.68 4340.34 30.12
RegCM4-6 CNRM-CM5 1589.19 3113.20 21.66
RegCM4-6 MPI-ESM-LR 1482.80 2988.44 20.82
GCM unseen during training (extrapolation case)
Appendix C Details on methods
C.1 Conditional independence assumptions for multi-step approach
@@
C.2 Conditional energy score loss
We call this the EnScale-loss. As before, we call the prediction error (first term) $\text{ES}_{\text{pred}}$ and the variability (second term) $\text{ES}_{\text{var}}$. This loss is minimized if and only if, for each predictor $x$, the predicted conditional distributions $q_{Y|X=x}$ indeed match the target conditionals $p_{Y|X=x}$. More formally, the score is minimized if and only if $q_{Y|X=x}=p_{Y|X=x}$ $p_{X}$-almost everywhere [46]. As for the unconditional loss in Eq. (1), if $p_{Y|X=x}=q_{Y|X=x}$ for each $x$, it holds that $\text{ES}_{\text{pred}}=\text{ES}_{\text{var}}$. During training, we use minibatching. We describe the standard batch-wise approximation of the loss in App. C.3, where we use two samples by the generative model for each input $x$. Each batch contains data from multiple GCM-RCM pairs, such that the loss is averaged across pairs.
C.3 Batch-wise approximation
The theoretical expected values (e.g. Eq. (6)) are impossible to calculate. However, during training, they can be replaced by averages over mini-batches of data, yielding an unbiased estimate of the expectation. Let $(x_{i},y_{i})\,i\in\{1,...,n\}$ be a mini-batch of data, and $\hat{y_{i}},\hat{y_{i}}^{\prime}\sim Q_{Y|X}(.|x_{i})$ two samples from the conditional distribution of the RCM data given the GCM input $x_{i}$. Then the finite-sample approximation of the loss is
@@ C.3 Batch-wise approximation
C.4 Model architecture
C.4.1 Coarse model
The first coarse model is a dense multilayer perceptron (MLP) and its architecture is as in [60]. To enable the model to capture difference between GCM-RCM pairs, we integrate the index for the RCM considered: We one-hot-encode the RCM index and append it to the input. In addition, the MLP is modified to include random noise that allows for stochastic output. As explained above, the generative model needs to take random initialization as input. Instead of just adding this noise to the input at the beginning, we add the randomness at each layer. We concatenate the hidden units with standard Gaussian noise, where the noise is scaled by a learnable parameter. This concatenation is passed into the next layer. Adding noise in multiple hidden layers allows to learn more complex representations of the Gaussian noise. For example, we find that this yields spatially correlated variability in the output.
@@ C.4.1 Coarse model
C.4.2 Super-resolution model
@@
As for the coarse model, we enable learning RCM-specific details by integrating the index for the RCM into the sparse local layers (see Sec. 2.4). In the first linear upsampling step, the learnable weights are RCM-specific, i.e. the model can learn a different set of weights for each RCM identifier. The stochastic refinement does not depend on the RCM and, thus, learned stochasticity is RCM-agnostic. In earlier stages of this work, we also tested learning a separate set of weights per GCM-RCM pair in both the first deterministic as well as the stochastic part, but found that using a joint model for the stochastic part reaches the same loss, while being more efficient computationally.
For EnScale-t, conditioning on the previous timestep is again integrated only in the first linear upsampling step: For each variable, we interpolate the low-resolution input to the next resolution, i.e., we increase resolution by a factor of 2 in each dimension (with nearest neighbor interpolation). We then concatenate the high-resolution prediction from the previous day as a separate channel and use both channels as an input for the linear upsampling.
C.5 Training
@@
C.6 Details on benchmarks
C.6.1 Deterministic neural network
We use a similar model architecture to the dense models in our generative approach (see Sec. C.4), using a fully-connected neural network with alternating linear maps and ReLU activation. However, unlike for the generative models, there is no noise input needed for the deterministic benchmark. The layers have sizes 3612, 500, 500, 500, 1024, 4096, 16384.
@@ C.6.1 Deterministic neural network
C.6.2 EasyUQ
EasyUQ’s predictions are optimal under a monotonicity assumption, in the sense that they minimize the CRPS. The monotonicity assumption describes that the targets $Y$ tend to increase as the regression model’s predictions $\hat{\mu}$ also increase. This is a natural requirement for a useful statistical model. However, if the regression model is deficient, this assumption might be violated. EasyUQ will still yield predictive distributions, but optimality is not guaranteed in this case. For more details on EasyUQ and its mathematical guarantees, see [72] and [26].
We implement EasyUQ based on the Python package isodisreg [73]. We train EasyUQ on the training data that was also used to train the NN-det, separately for each location. After training, we retrieve the predicted distributions on the test data and sample from them, independently in each location, time point and for each variable. For this, we slightly adjust the default prediction method from the Python package in order to be able to take in a different set of probabilities for each point in time. These samples can then be compared to the samples from EnScale.
C.7 Analogues
@@
C.8 Proof of variance decomposition
@@
We explained above (see App. C.1) that our approach to downscaling via coarse correction assumes $Y\mathrel{\perp\!\!\!\perp}X|Z$. Hence, $\operatorname{Var}^{\iota}(Y|Z,X)=\operatorname{Var}^{\iota}(Y|Z)$ and $E^{\iota}[Y|Z,X]=E^{\iota}[Y|Z]$, which concludes the proof.